# Robust and Computation-Aware Gaussian Processes

### v1

@@ 3.3 Joint model parameters and design selection strategy
To fully take advantage of RCaGP in BO, we leverage the expected lower bound utility (EULBO) framework introduced by [28]. At acquisition time, instead of optimizing the surrogate’s inducing points, variational parameters, and hyperparameters, only then to find a design maximizing the acquisition function, a global criterion is maximized in an end-to-end manner:
@@
where $y_{t}^{\ast}$ denotes the best evaluation observed so far and $\mathrm{softplus}:{\bf{x}}\mapsto\log(1+\exp({\bf{x}}))$. This reformulation replaces the commonly used ReLU function with a softplus function, guaranteeing the utility function remains strictly positive whenever $f({\bf{x}})\geq y^{\ast}$. Moreover, EULBO can be extended to support batch BO using Monte Carlo batch mode [4, 47]. Given a set of candidates ${\bf{X}}=\{{\bf{x}}_{1},\dots,{\bf{x}}_{q}\}$, the expected utility function corresponding to the $q-$improvement utility is:
@@
For efficiency, we impose the sparse block structure on the action matrices ${\bf{S}}_{i}\in^{n\times i}$, following [46]. Specifically, we enforce each block to be column vector ${\bf{s}}_{j}\in^{k\times 1}$, with $k=n/i$ entries, so that the number of trainable parameters is $k\times i=n$ (training data size). It is worth noticing that unlike EULBO with SVGP, which maximizes variational parameters and inducing locations, our approach directly optimizes the action matrices ${\bf{S}}_{i}$.
Maximizing the ELBO independently of the posterior-expected utility function can result in suboptimal data acquisition decisions, as the ELBO is primarily designed to model observed data [27]. In contrast, EULBO considers how the surrogate performs when selecting the next query, effectively guiding the solution of ELBO optimization towards high utility regions [28]. Our RCaGP model further balances such aggressive behavior by incorporating the combined uncertainty, preventing overconfidence during candidate query selection. Moreover, RCaGP enhances robustness in scenarios where function evaluations occasionally produce outliers.
@@ 4 Theoretical Analysis
This section presents a theoretical analysis of RCaGP. Proposition 4.1 establishes its robustness to outliers, while Proposition 4.2 links its uncertainty estimates to worst-case error, ensuring conservative predictions, thus addressing the challenges outlined in Section 1.
We demonstrate the robustness of RCaGP through the posterior influence function (PIF), similarly to [2, Proposition 3.2]. The PIF is a criterion proving robustness to misspecification in observation error [11, 26]. For this purpose, we define the dataset ${\mathcal{D}}=\{({\bf{x}}_{j},y_{j})\}_{j=1}^{n}$ and the corresponding contaminated dataset ${\mathcal{D}}_{m}^{c}=({\mathcal{D}}\setminus({\bf{x}}_{m},y_{m}))\cup({\bf{x}}_{m},y_{m}^{c})$, indexed by $m\in\{1,\dots,n\}$. The impact of $y_{m}^{c}$ on inference is measured through PIF, expressed as Kullback-Leibler (KL) divergence between RCaGP’s contaminated posterior $p({\bf{f}}|{\mathcal{D}}_{m}^{c})$ and its uncontaminated counterpart $p({\bf{f}}|{\mathcal{D}})$:
@@
PIF views the KL-divergence as a function of $|y_{m}^{c}-y_{m}|$. In principle, we can consider any divergence that is not uniformly bounded. Here, we choose KL-divergence since it provides a closed-form expression for multivariate Gaussian. A posterior is robust if $\sup_{y\in\mathcal{Y}}|\mathrm{PIF}(y_{m}^{c},{\mathcal{D}})|<\infty$ [2]. We then show that RCaGP’s PIF is bounded through the following proposition:
@@
The constraint $\sup_{{\bf{x}},y}w({\bf{x}},y)<\infty$ ensures no observation has infinite weight, and $\sup_{{\bf{x}},y}y,w^{2}({\bf{x}},y)<\infty$ guarantees that $w$ down-weights observations at least at rate $1/y$. The proof sketch shows that each term in Equation 18 is bounded. We leverage the pseudo-inverse property of positive semi-definite matrices and matrix norm bounds on the low-rank approximation to relate RCGP and CaGP. The full proof is in Appendix D. Crucially, the PIF of CaGP is unbounded, indicating a lack of robustness to outliers (see Section F.1).
Next, we show that RCaGP’s uncertainty captures the worst-case error over all latent functions, paralleling [45, Theorem 2]. We assume the difference between the latent function $g$ and the shrinkage term $m_{w}$ lies in an RKHS. Under this assumption, we establish the link between RCaGP’s uncertainty estimates and the worst-case error through the following proposition:
@@
The first equation in Proposition 4.2 shows that RCaGP’s uncertainty $\hat{k}_{i}(.,.)$ captures the worst-case error between the latent function $g$ and the posterior mean $\hat{\mu}_{i}(.)$. The second shows that the computational uncertainty $\hat{\sigma}_{i}^{2}$ captures the worst-case error between the RCGP and RCaGP posterior means, induced by approximation actions such as inducing points. This added uncertainty prevents overconfidence when modeling the latent function. The proof follows [45], leveraging [Proposition 3.9] of [20] and the RKHS associated with RCGP. Full details are in Appendix E. Lastly, the convergence of RCaGP in terms of mean function in RKHS norm is established in Section F.2.
@@ 5 Expert-guided robust mean prior definition
Our proposed RCaGP enforces outlier robustness through the weight function $w$, which—as discussed in Section 3.1 and Proposition 4.2—critically depends on the prior mean $m$. Motivated by these insights, we propose defining a robust and informative mean prior using domain expert feedback.
@@
This corresponds to averaging over the $\hat{n}$ nearest identified outliers, weighted by a kernel $l$ and the expected confidence in each correction. In practice, we average over all identified outliers and use only posterior expectations, prioritizing computational efficiency over full Bayesian inference. The full generative process and toy illustration appear in Figure 2 (left and middle). Additional inference details are provided in Appendix G.
@@ 6 Related Work
Outlier-robust Gaussian processes fall into three main categories. The first replaces Gaussian noise with heavy-tailed alternatives—e.g., Student-$t$[19, 34], Huber[1], Laplace [22], data-dependent [12], or mixture models [7, 24]. These approaches break conjugacy and require approximate inference. The second class removes suspected outliers before applying standard inference [23, 31, 3], but this is often unreliable and computationally expensive in high dimensions. The third approach, RCGP [2], builds on generalized Bayesian inference [6, 18, 21] to down-weight outliers without discarding data. We adopt RCGP as the foundation for RCaGP due to its strong performance in regression with outliers, its preserved conjugacy, and greater computational efficiency—enabling more accurate posterior inference and better predictive performance.
SVGPs are the standard approach for scaling Gaussian processes to large-scale Bayesian optimization [13, 37, 43, 42]. [30] improved inducing point selection by optimizing their weighted Gram matrix via a determinantal point process. [28] formulated acquisition as expected utility, enabling joint optimization of queries, variational parameters, hyperparameters, and inducing points in an end-to-end fashion. [25] proposed an online conditioning strategy that updates SVGPs with new data without reoptimizing variational parameters, enabling look-ahead acquisition in BO.
Computation-aware Gaussian processes are a novel class of GPs that accounts for computational uncertainty, due to approximations in matrix-vector multiplications [45]. These methods fall under the domain of probabilistic numerics, which quantify additional uncertainties arising from approximation errors [32, 39, 33]. In that respect, [46] introduced a batch update algorithm for CaGP and proposed an ELBO objective to facilitate model selection. [15] extended CaGPs by proposing a calibrated probabilistic linear solver, resulting in a calibrated computation-aware GP model.
@@ 7 Experimental results
We evaluate the performance of our proposed RCaGP against concurrent baselines on a range of regression datasets, followed by several high-throughput Bayesian Optimization tasks. We conclude by presenting insights gained after carrying out dedicated ablation studies.
GP baselines. We compare RCaGP against RCSVGP [2], SVGP [16], and CaGP [46]. Finally, we set the number of actions for RCaGP and CaGP, and the number of inducing points for SVGP and RCSVGP, to $25$. All baselines are equipped with a Matern $5/2$ kernel and were implemented using GPyTorch [9]. Model hyperparameters are optimized using the ELBO (Equation 13).
Outlier contamination protocol. In our experiments, we follow the settings described by [2]. Unless stated otherwise in dedicated ablation studies, for regression datasets, we sample uniformly at random 10% of the training dataset input-output pairs $({\bf{x}}_{i},y_{i})$, and replace the $y_{i}$’s by asymmetric outliers, i.e., via subtraction of noise sampled from a uniform distribution $\mathcal{U}(3\bar{\sigma},9\bar{\sigma})$, with $\bar{\sigma}$ being the standard deviation of the original observations. For BO tasks, while running the optimization loop, each evaluation has a $25\%$ chance of returning an asymmetric outlier, obtained by addition of noise sampled from a uniform distribution $\mathcal{U}(1\bar{\sigma},2\bar{\sigma})$, with $\bar{\sigma}$ being is the estimated standard deviation of the normalized evaluation function. Finally, for the weight function $w$ (Equation 12), we set the soft threshold $c=Q_{n}(|\mathbf{y}-{\bf{m}}|,1-\epsilon)$ with $\epsilon=0.2$, with $Q_{n}$ the $(1-\epsilon)$-quantile of $|\mathbf{y}-{\bf{m}}|$.
Bayesian Optimization settings. As described during Section 3.3, we integrate the baselines into EULBO, building on the implementation of [28]. We initialize the optimization process for all baselines with $250$ data points sampled uniformly across the search space. All surrogates employ expected improvement (EI) as the acquisition function. Standard BO is applied across all tasks, while trust-region BO (TuRBO [8]) is additionally employed for high-dimensional tasks.
@@ 7.1 Regression on UCI datasets
We evaluate GP regression on four UCI datasets with asymmetric outliers (details in Section H.1); results are shown in the left part of Table 1. RCaGP consistently delivers the best performance, combining strong predictive accuracy and well-calibrated uncertainty with acceptable runtimes. Notably, it significantly outperforms both CaGP and RCSVGP in this setting, indicating that the integration of robustness and approximation-awareness is effective.
To test robustness across contamination types, we further consider two additional outlier scenarios, uniform and focused outliers, inspired by [2] (definitions in Section H.3). The corresponding results are shown in Table S1 and Table S2. RCaGP remains the top performer overall, particularly in uniform outliers for MAE, and focused outliers for NLL. These findings support its robustness across contamination types. More importantly, because outlier presence is not always known in advance, we also evaluate performance under clean conditions (right side of Table 1). RCaGP remains competitive and often superior—even without injected noise. One exception is the Energy dataset, where CaGP marginally outperforms RCaGP. This suggests that even in the absence of outliers, the weight function $w$ used by RCaGP (Equation 12) proves useful. An alternative explanation would be that these datasets already contain outliers from the start.
Overall, these results demonstrate that RCaGP is more than the sum of its parts: the combination of RCGP and CaGP is not merely additive, but synergistic, consistently outperforming either component alone—both in outlier-rich and clean regimes.

### v2

@@ 3.3 Joint model parameters and design selection strategy
To fully take advantage of RCaGP in BO, we leverage the expected lower bound utility (EULBO) framework introduced by [maus2024approximation]. At acquisition time, instead of optimizing the surrogate’s inducing points, variational parameters, and hyperparameters, only then to find a design maximizing the acquisition function, a global criterion is maximized in an end-to-end manner:
@@
where $y_{t}^{\ast}$ denotes the best evaluation observed so far and $\mathrm{softplus}:{\bf{x}}\mapsto\log(1+\exp({\bf{x}}))$. This reformulation replaces the commonly used ReLU function with a softplus function, guaranteeing the utility function remains strictly positive whenever $f({\bf{x}})\geq y^{\ast}$. Moreover, EULBO can be extended to support batch BO using Monte Carlo batch mode [balandat2020botorch, wilson2018maximizing]. Given a set of candidates ${\bf{X}}=\{{\bf{x}}_{1},\dots,{\bf{x}}_{q}\}$, the expected utility function corresponding to the $q$-improvement utility is:
@@
For efficiency, we impose the sparse block structure on the action matrices ${\bf{S}}_{i}\in^{n\times i}$, following [wenger2024computation]. Specifically, we enforce each block to be column vector ${\bf{s}}_{j}\in^{k\times 1}$, with $k=n/i$ entries, so that the number of trainable parameters is $k\times i=n$ (training data size). It is worth noticing that unlike EULBO with SVGP, which maximizes variational parameters and inducing locations, our approach directly optimizes the action matrices ${\bf{S}}_{i}$.
Maximizing the ELBO independently of the posterior-expected utility function can result in suboptimal data acquisition decisions, as the ELBO is primarily designed to model observed data [matthews2016sparse]. In contrast, EULBO considers how the surrogate performs when selecting the next query, effectively guiding the solution of ELBO optimization towards high utility regions [maus2024approximation]. Our RCaGP model further balances such aggressive behavior by incorporating the combined uncertainty, preventing overconfidence during candidate query selection. Moreover, RCaGP enhances robustness in scenarios where function evaluations occasionally produce outliers.
@@ 4 Theoretical Analysis
This section presents a theoretical analysis of RCaGP. Proposition˜4.1 establishes its robustness to outliers, while Proposition˜4.2 links its uncertainty estimates to worst-case error, ensuring conservative predictions, thus addressing the challenges outlined in Section˜1.
We demonstrate the robustness of RCaGP through the posterior influence function (PIF), similarly to [altamirano2023robust, Proposition 3.2]. The PIF is a criterion proving robustness to misspecification in observation error [ghosh2016robust, matsubara2022robust]. For this purpose, we define the dataset ${\mathcal{D}}=\{({\bf{x}}_{j},y_{j})\}_{j=1}^{n}$ and the corresponding contaminated dataset ${\mathcal{D}}_{m}^{c}=({\mathcal{D}}\setminus({\bf{x}}_{m},y_{m}))\cup({\bf{x}}_{m},y_{m}^{c})$, indexed by $m\in\{1,\dots,n\}$. The impact of $y_{m}^{c}$ on inference is measured through PIF, expressed as Kullback-Leibler (KL) divergence between RCaGP’s contaminated posterior $p({\bf{f}}|{\mathcal{D}}_{m}^{c})$ and its uncontaminated counterpart $p({\bf{f}}|{\mathcal{D}})$:
@@
PIF views the KL-divergence as a function of $|y_{m}^{c}-y_{m}|$. In principle, we can consider any divergence that is not uniformly bounded. Here, we choose KL-divergence since it provides a closed-form expression for multivariate Gaussian. A posterior is robust if $\sup_{y\in\mathcal{Y}}|\mathrm{PIF}(y_{m}^{c},{\mathcal{D}})|<\infty$ [altamirano2023robust]. We then show that RCaGP’s PIF is bounded through the following proposition:
@@
The constraint $\sup_{{\bf{x}},y}w({\bf{x}},y)<\infty$ ensures no observation has infinite weight, and $\sup_{{\bf{x}},y}y,w^{2}({\bf{x}},y)<\infty$ guarantees that $w$ down-weights observations at least at rate $1/y$. The proof sketch shows that each term in Equation˜18 is bounded. We leverage the pseudo-inverse property of positive semi-definite matrices and matrix norm bounds on the low-rank approximation to relate RCGP and CaGP. The full proof is in Appendix˜D. Crucially, the PIF of CaGP is unbounded, indicating a lack of robustness to outliers (see Section˜F.1).
Next, we show that RCaGP’s uncertainty captures the worst-case error over all latent functions, paralleling [wenger2022posterior, Theorem 2]. We assume the difference between the latent function $g$ and the shrinkage term $m_{w}$ lies in an RKHS. This holds when $\|\nabla_{y}\log w^{2}({\bf{x}},y)\|^{2}_{{\mathcal{H}}_{k}}<\infty$. In the case of our robust weight function, the correction term is a nonlinear rational function. Generally, such functions do not lie in the RKHS associated with common kernels (e.g., RBF, Matérn), unless specific and uncommon conditions are met—such as the RKHS being closed under composition with the given nonlinearity. A practical approach to ensure it belongs to the RKHS is to project the correction term onto the RKHS. Alternatively, one could use or design a kernel whose RKHS explicitly accommodates this class of nonlinear transformations. Under this assumption, we establish the link between RCaGP’s uncertainty estimates and the worst-case error through the following proposition:
@@
The first equation in Proposition˜4.2 shows that RCaGP’s uncertainty $\hat{k}_{i}(.,.)$ captures the worst-case error between the latent function $g$ and the posterior mean $\hat{\mu}_{i}(.)$. Instead of pretending the model is fully correct, we allow a structured form of misspecification, namely, the part captured by ${\bf{m}}_{w}$. Proposition˜4.2 then guarantees that any remaining discrepancy is still bounded by our variance. Even when classical Bayesian inference may fail to offer meaningful uncertainty due to model misspecification or outliers, RCaGP maintains a rigorous worst-case interpretation of its predictive variance, so long as the residual is sufficiently regular. This mirrors the CaGP assumption in spirit: just as CaGP analyzes worst-case error over functions in the RKHS, RCaGP does the same for the residual, which can be viewed as the portion of the signal that the robust update leaves uncorrected. The second shows that the computational uncertainty $\hat{\sigma}_{i}^{2}$ captures the worst-case error between the RCGP and RCaGP posterior means, induced by approximation actions such as inducing points. The proof follows [wenger2022posterior], leveraging [kanagawa2018gaussian, Proposition 3.9] and the RKHS associated with RCGP. Full details are in Appendix˜E. Lastly, the convergence of RCaGP in terms of mean function in RKHS norm is established in Section˜F.2.
@@ 5 Expert-guided robust mean prior definition
Our proposed RCaGP enforces outlier robustness through the weight function $w$, which—as discussed in Section 3.1 and Proposition˜4.2—critically depends on the prior mean $m$. Motivated by these insights, we propose defining a robust and informative mean prior using domain expert feedback.
@@
This corresponds to averaging over the $\hat{n}$ nearest identified outliers, weighted by a kernel $l$ and the expected confidence in each correction. In practice, we average over all identified outliers and use only posterior expectations, prioritizing computational efficiency over full Bayesian inference. The full generative process and toy illustration appear in Figure˜2 (left and middle). Additional inference details are provided in Appendix˜G.
@@ 6 Related Work
Outlier-robust Gaussian processes fall into three main categories. The first replaces Gaussian noise with heavy-tailed alternatives—e.g., Student-tt[jylanki2011robust, ranjan2016robust], Huber[algikar2023robust], Laplace [kuss2006gaussian], data-dependent [goldberg1997regression], or mixture models [daemi2019identification, lu2023robust]. These approaches break conjugacy and require approximate inference. The second class removes suspected outliers before applying standard inference [li2021robust, park2022robust, andrade2023robust], but this is often unreliable and computationally expensive in high dimensions. The third approach, RCGP [altamirano2023robust], builds on generalized Bayesian inference [bissiri2016general, jewson2018principles, knoblauch2022optimization] to down-weight outliers without discarding data. We adopt RCGP as the foundation for RCaGP due to its strong performance in regression with outliers, its preserved conjugacy, and greater computational efficiency—enabling more accurate posterior inference and better predictive performance.
SVGPs are the standard approach for scaling Gaussian processes to large-scale Bayesian optimization [griffiths2020constrained, stanton2022accelerating, vakili2021scalable, tripp2020sample]. [moss2023inducing] improved inducing point selection by optimizing their weighted Gram matrix via a determinantal point process. [maus2024approximation] formulated acquisition as expected utility, enabling joint optimization of queries, variational parameters, hyperparameters, and inducing points in an end-to-end fashion. [maddox2021conditioning] proposed an online conditioning strategy that updates SVGPs with new data without reoptimizing variational parameters, enabling look-ahead acquisition in BO.
Computation-aware Gaussian processes are a novel class of GPs that accounts for computational uncertainty, due to approximations in matrix-vector multiplications [wenger2022posterior]. These methods fall under the domain of probabilistic numerics, which quantify additional uncertainties arising from approximation errors [pfortner2022physics, tatzel2023accelerating, pfortner2024computation]. In that respect, [wenger2024computation] introduced a batch update algorithm for CaGP and proposed an ELBO objective to facilitate model selection. [hegde2024calibrated] extended CaGPs by proposing a calibrated probabilistic linear solver, resulting in a calibrated computation-aware GP model.
@@ 7 Experimental results
We evaluate the performance of our proposed RCaGP against concurrent baselines on a range of regression datasets, including a real-world dataset, followed by several high-throughput Bayesian Optimization tasks. We conclude by presenting insights gained after carrying out dedicated ablation studies. Our implementation is available at https://github.com/MarshalArijona/RCaGP.
GP baselines. We compare RCaGP against several baseline methods: RCSVGP [altamirano2023robust], SVGP [hensman2013gaussian], and CaGP [wenger2024computation]. For the UCI regression benchmarks, we further include comparisons with SVGP models using a Student-$t$ likelihood [jylanki2011robust] and SVGP with relevance pursuit (RRP), adapted from [ament2024robust] and formally derived in Section F.3. We set the number of actions (for RCaGP and CaGP) or inducing points (for SVGPs and RCSVGP) to 25. All baselines employ a Matérn-$5/2$ kernel and are implemented in GPyTorch [gardner2018gpytorch]. Model hyperparameters are optimized by maximizing the ELBO (Equation 13).
Outlier contamination protocol. In our experiments, we follow the settings described by [altamirano2023robust]. Unless stated otherwise in dedicated ablation studies, for regression datasets, we sample uniformly at random 10% of the training dataset input-output pairs $({\bf{x}}_{i},y_{i})$, and replace the $y_{i}$’s by asymmetric outliers, i.e., via subtraction of noise sampled from a uniform distribution $\mathcal{U}(3\bar{\sigma},9\bar{\sigma})$, with $\bar{\sigma}$ being the standard deviation of the original observations. For BO tasks, while running the optimization loop, each evaluation has a $25\%$ chance of returning an asymmetric outlier, obtained by addition of noise sampled from a uniform distribution $\mathcal{U}(1\bar{\sigma},2\bar{\sigma})$, with $\bar{\sigma}$ being is the estimated standard deviation of the normalized evaluation function. Finally, for the weight function $w$ (Equation 12), we set the soft threshold $c=Q_{n}(|\mathbf{y}-{\bf{m}}|,1-\epsilon)$ with $\epsilon=0.2$, with $Q_{n}$ the $(1-\epsilon)$-quantile of $|\mathbf{y}-{\bf{m}}|$.
Bayesian Optimization settings. As described during Section 3.3, we integrate the baselines into EULBO, building on the implementation of [maus2024approximation]. We initialize the optimization process for all baselines with $250$ data points sampled uniformly across the search space. All surrogates employ expected improvement (EI) as the acquisition function. Standard BO is applied across all tasks, while trust-region BO (TuRBO [eriksson2019scalable]) is additionally employed for high-dimensional tasks.
@@ 7.1 Regression on UCI datasets
We evaluate GP regression on four UCI datasets with asymmetric outliers (details in Section H.1); results are shown in the left part of Table 1. RCaGP delivers the overall best performance, combining strong predictive accuracy and well-calibrated uncertainty with acceptable runtimes, and being only surpassed in the Yacht dataset by RRP, a much slower baseline. Notably, it significantly outperforms both CaGP and RCSVGP in this setting, indicating that the integration of robustness and approximation-awareness is effective.
To test robustness across contamination types, we further consider two additional outlier scenarios, uniform and focused outliers, inspired by [altamirano2023robust] (definitions in Section H.3). The corresponding results are shown in Table S1 and Table S2. RCaGP remains the top performer overall, particularly in uniform outliers for MAE, and focused outliers for NLL. These findings support its robustness across contamination types. More importantly, because outlier presence is not always known in advance, we also evaluate performance in the absence of outliers (right side of Table 1). RCaGP remains superior, except for the Energy dataset, where CaGP marginally outperforms RCaGP. This suggests that even in the absence of outliers, the weight function $w$ used by RCaGP (Equation 12) proves useful. An alternative explanation would be that these datasets already contain outliers from the start.
Overall, these results demonstrate that RCaGP is more than the sum of its parts: the combination of RCGP and CaGP is not merely additive, but synergistic, consistently outperforming either component alone, both in outlier-rich and clean regimes.
@@
Student-t MAE 0.962 ± 0.037 1.015 ± 0.028 0.907 ± 0.061 1.025 ± 0.079 0.896 ± 0.041 0.970 ± 0.025 0.815 ± 0.071 0.965 ± 0.074
NLL 1.619 ± 0.039 1.606 ± 0.024 1.564 ± 0.067 1.658 ± 0.085 1.593 ± 0.063 1.572 ± 0.026 1.559 ± 0.100 1.627 ± 0.092
clock-time 1.210 ± 0.240 1.800 ± 0.400 1.350 ± 0.260 0.990 ± 0.090 1.330 ± 0.050 1.330 ± 0.080 0.890 ± 0.150 0.730 ± 0.020
RRP MAE 0.550 ± 0.050 0.597 ± 0.112 0.356 ± 0.051 0.644 ± 0.106 0.815 ± 0.103 0.775 ± 0.118 0.708 ± 0.035 1.052 ± 0.100
NLL 1.147 ± 0.050 1.240 ± 0.117 0.896 ± 0.030 1.306 ± 0.121 1.527 ± 0.194 1.547 ± 0.214 1.179 ± 0.045 / 1.884 ± 0.179
clock-time 5.120 ± 0.100 3.840 ± 0.090 7.200 ± 0.160 3.010 ± 0.080 5.260 ± 0.120 3.730 ± 0.100 6.980 ± 0.110 3.010 ± 0.060
7.2 Twitter Flash Crash
We illustrate the real-world applicability of RCaGP by evaluating its performance and other baselines on the Dow Jones Industrial Average (DJIA) index for April 17, 2013. On this day, the Associated Press Twitter account was hacked, posting a false report of an explosion at the White House and an injury to the U.S. President. The resulting panic triggered a sudden market sell-off, followed by a rapid rebound, creating a brief interval during which the DJIA did not reflect the true economic state of the U.S. stock market. This scenario provides a unique stress test for evaluating model robustness under sudden, anomalous market conditions. Observations during this period can reasonably be treated as outliers. Figure˜3 (left) shows the DJIA index alongside the predictive means of RCaGP and other baselines. The predictive means of SVGP, RCSVGP, and CaGP exhibit noticeable deviations before or after 1 PM, while RCaGP yields a more reliable estimate, with only minor discrepancies around 2 PM. These results suggest that insufficient robustness or computational uncertainty in the baseline models contributes to their degraded performance under sudden market shocks. In the right panel of Figure˜3, we plot the weight functions of RCaGP and RCSVGP. Although their weight functions exhibit similar profiles, they yield markedly different mean predictions. This supports our claim that robustness alone, without an appropriate treatment of computational uncertainty, can degrade the performance of approximate Gaussian process models.
Figure 3: Left: Predictive mean of various approximate GP models on the DJIA index with a constant mean prior. Outliers affect the baselines, whereas RCaGP maintains robustness. Right: learned weight functions $w({\bf{x}},y)$ for RCaGP and RCSVGP, illustrating their influence on the regression.